A ferry leaves a harbor and travels to a resort island at an average speed of 22 mph. On the return trip, because of fog, the ferry travels at an average speed of 10 mph. The total time for the trip is 4 h. How far is the island from the harbor?
Answer by macston(5194)

Distance one way =D
D/22+D/10=4 find common denominator
D/22(5/5)+D/10(11/11)=4 hrs
5D/110+11D/110=4 hrs
16D/110=4 hrs multiply each side by 110
16D=440 Divide each side by 16
D=27.5 mi ANSWER The island is 27.5 mi from the harbor
CHECK
round trip=time outbound(22 mph)+time inbound(10 mph)
4 hrs=27.5mi/22mph+ 27.5mi/10mph
4 hrs=1.25 hrs+2.75 hrs
4 hrs=4 hrs